Assume: A 78 g basketball is launched at an
angle of 41.9◦ and a distance of 18.5 m from
the basketball goal. The ball is released at the
same height (ten feet) as the basketball goal’s
height.
A basketball player tries to make a long
jump-shot as described above.
The acceleration of gravity is 9.8 m/s2 .
What speed must the player give the ball?
Answer in units of m/s.

Answer by stanbon(75887)

Assume: A 78 g basketball is launched at an
angle of 41.9◦ and a distance of 18.5 m from
the basketball goal. The ball is released at the
same height (ten feet) as the basketball goal’s
height.
A basketball player tries to make a long
jump-shot as described above.
The acceleration of gravity is 9.8 m/s2 .
What speed must the player give the ball?
Answer in units of m/s.
------------------------------------------------
Plot the points (0,0) and (18.5,0) on the x-axis
(0,0) is the point of release; (18.5,0) is the basket
----------
Equation: s(t) = -9.8t^2 + vt
s is the height at time t
v is the release velocity
-----------------------------
Your Problem:
-9.8t^2 + vt = 0
t(-9.8t + v) = 0
t = 0 or t = v/9.8
-------------------------
To solve for "v" you need to know the time from release to basket.
Your post refers to a "jump-shot as described above". Does that
description include the time? If not you cannot solve for "v".
===================================================================
Cheers,
Stan H.